OpenStudy (anonymous):

the expression 10/(3 + i) is equivalent to: 1) 3 - i 2) (15 + 5i)/4 3) 3 + i 4) 5/4

OpenStudy (anonymous):

=10/(3+i) * (3-i)/(3-i) =10(3-i)/10 =3-i

OpenStudy (anonymous):

3 - i

OpenStudy (anonymous):

3-i

OpenStudy (anonymous):

3-1